### v1

# Improving the mean-field approximation in continuous models of population dynamics with nonlocal dispersal: applications to vegetation pattern formation

Abstract
Spot patterns, in which vegetation patches form a hexagonal lattice, are frequent in nature and could serve as an early-warning indicator of abrupt vegetation collapses. Consequently, they have been intensively studied using both individual-based models and density-based field equations. Yet, the relationship between these two approaches remains unclear, particularly in scenarios where vegetation dynamics exhibit strong long-range spatial correlations and traditional mean-field approximations fail. To solve this issue, we develop a new method that refines mean-field approximations by describing both the dynamics of the biomass density field and its spatial correlations. This new approach harnesses the strengths of both individual and density-based mdoels, treating spatial correlations explicitly and allowing for the identification of spatial instabilities resulting in periodic patterns. Our results indicate that this new approximation predicts the parameter regimes where regular periodic patterns emerge more accurately than mean-field models, suggesting that it could provide a more robust framework to perform further nonlinear analysis.
1 Introduction
Populations of sessile organisms exhibit diverse spatial patterns, from randomly scattered individuals to aggregates that are periodically distributed in space [1, 2]. The latter are particularly interesting from a modeling point of view because they raise questions about how interactions operating at the scale of organisms lead to organized structures at much larger spatial scales [3, 4, 5]. Moreover, regular patterns could enhance population productivity, resilience, and resistance to higher environmental stresses, underscoring their ecological significance [6, 7, 8, 9].
Plants in water-limited ecosystems provide a good system to investigate regular pattern formation in populations of sessile and semi-sessile organisms [10]. Several studies using aerial and satellite images from different locations have identified a variety of patterns in these systems, including gaps, stripes, rings, labyrinths, and spots [11, 12, 13, 14]. These patterns are primarily shaped by environmental factors such as gradients in resource availability, interactions between plants and other organisms, and local topography [15, 16, 17, 18, 14, 19]. For example, along a gradient of increasing aridity, vegetation pattern shapes shift from gaps to labyrinths and ultimately to spots [12]. Therefore, these spot patterns, where approximately circular patches of vegetation form a hexagonal lattice in a matrix of bare soil [20], could indicate proximity to non-vegetated states [8]. This potential role as an early-warning indicator of desertification processes makes spot patterns ecologically relevant. However, the long timescales governing vegetation dynamics and the limited availability of long-term datasets pose challenges for empirically investigating the ecological significance of vegetation patterns [but see 21, 22]. Alternatively, researchers have used spatially explicit theoretical models to explain both how spot patterns could form and the types of desertification processes they might precede [23, 24, 25, 26, 27]. Broadly, existing models can be classified into two categories: discrete individual-based (IBM), and continuous density-based models.
IBMs describe a population as a spatial distribution of point-like organisms that reproduce and die according to probabilistic rules that encode different interactions and dispersal mechanisms [28, 29]. These models are built at the level of individual plants, which makes it possible to study spatial patterns resolving the patch structure, and are easy to simulate computationally. However, the computational overhead associated with these simulations scales rapidly with the system size and the complexity of interactions, and analytical techniques to study them are limited to providing predictions of population size and its spatial correlations at equilibrium [30, 31, 32]. In continuous models, on the other hand, the vegetation biomass is a field that changes in space and time following a differential equation [33, 34]. This approach is easier to treat mathematically, which makes it possible to predict the dynamics of certain spatial processes, such as the direction of ‘desertification’ fronts [35]. However, density-based models do not resolve individual plants, which limits how much they inform about patch composition. This limitation turns them inappropriate for studying random and irregular aggregated patterns [36] unless they consider environmental heterogeneity [37, 38, 39, 40] or demographic fluctuations explicitly [41, 42, 43, 44].
Continuous models based on differential equations and IBMs are, thus, two complementary ways of investigating vegetation patterns and their ecological consequences, each presenting its limitations and strengths. Moreover, even though there are techniques to derive effective continuous models starting from individual-based dynamical rules [45], very few studies follow this approach, which makes it unclear how IBMs and continuous models based on field equations relate and compare. Here, we tackle this issue by analyzing a simple IBM with a minimal set of ingredients leading to regular spatial patterns. The existence of regular patterns makes this model tractable both using analytical approximations to the individual-based stochastic dynamics and a continuous equation for the biomass density field. Combining these two approaches, we show how a mean-field density equation fails in predicting the pattern formation regime, contrary to previous studies with motile organisms [46, 47], and use spatial moment dynamics equations to introduce a novel method that corrects this discrepancy. This corrected density equation could be useful for performing further nonlinear analyses that exploit the strengths of density-based models [48], such as determining transitions between pattern states, the existence of localized structure solutions, or the dynamics of fronts, to mention a few.
2 Model formulation and analytical approximations
2.1 Description of the stochastic dynamics
At the individual level, we model the population dynamics using a spatially explicit stochastic logistic model [36, 28]. We assume that organisms are identical except for their location and consider that they inhabit a two-dimensional landscape $\Omega$, with area $|\Omega|=A$. Therefore, the state of the population at a given time $t$ is fully determined by the set of plant locations ${\mathbf{x}}_{i}$ $[i=1,\ldots N(t)]$, where $N(t)$ is the number of plants at time $t$. Altogether, these locations define a spatial pattern $p(\mathbf{x},t)=\sum_{i}\delta(\mathbf{x}-\mathbf{x}_{i})$.
To keep our analysis as simple as possible, we consider a simple logistic-like birth-death dynamics in continuous time [36]. The $i$-th individual, located at $\mathbf{x}_{i}$, reproduces and disperses a new organism at location ${\mathbf{x}}$ at a rate
$B_{i}(\mathbf{x})=b\,\phi_{\mathrm{D}}({\mathbf{x}}-{\mathbf{x}}_{i}),$ (2.1)
where $\phi_{\mathrm{D}}({\mathbf{x}}-{\mathbf{x}}_{i})$ is a dispersal kernel that defines the probability that individuals born at location ${\mathbf{x}}_{i}$ establish at location ${\mathbf{x}}$. Because it acts as a probability density function, $\phi_{\mathrm{D}}$ must be normalized to unity. Several empirical studies have measured these dispersal kernels across species and investigated their importance in shaping ecological dynamics [49, 50]. Here, we follow a more common choice in the modeling literature and assume this kernel to be a Normal distribution centered at ${\mathbf{x}}_{i}$ and with variance $s$ [28, 32]. Conversely, each organism in a given spatial distribution $p$ dies at a rate
$D_{i}(p)=d+g\sum_{j\neq i}\phi_{\mathrm{C}}({\mathbf{x}}_{i}-{\mathbf{x}}_{j}),$ (2.2)
where the first term on the right side corresponds to baseline mortality, and the second term accounts for density-dependent effects due to intraspecific competition. $\phi_{\mathrm{C}}$ is the competition kernel that weighs the influence of an organism at ${\mathbf{x}}_{j}$ on the focal individual at ${\mathbf{x}}_{i}$. The shape of the competition kernel is a key ingredient in determining the type of spatial patterns that can form [46, 51]. In the simulations, we will use a Gaussian with mean zero and standard deviation $r_{c}$ and a top-hat competition kernel of the form
$\phi_{\mathrm{C}}({\mathbf{x}}_{i}-{\mathbf{x}}_{j})=\begin{dcases}\frac{1}{\pi r_{c}^{2}}&\mbox{if}\ \rvert{\mathbf{x}}_{i}-{\mathbf{x}}_{j}\rvert\leq r_{c},\\ 0&\mbox{otherwise}.\\ \end{dcases}$ (2.3)
In both cases, $r_{c}$ defines the characteristic spatial scale of the competition. We performed all IBM simulations in a square domain of lateral length $L=20$, using periodic boundary conditions and demographic rates $d/b=0.1$, with $b=3.05$ and $g=1$.
2.2 Spatial moment dynamics equations
The discrete dynamics in Section 2.1 can be approximated at the macroscale in terms of the moments of the population spatial pattern. For the kind of spatial point process considered in the IBM, the first spatial moment, $Z_{1}(t)$, is the average density of individuals. The second spatial moment, $Z_{2}(\bm{\xi},t)$, is defined as the average density of pairs of individuals separated by a displacement $\bm{\xi}$. Similarly, the third spatial moment, $Z_{3}(\bm{\xi},\bm{\xi}^{\prime},t)$, corresponds to the average density of triplet of individuals separated by displacements $\bm{\xi}$ and $\bm{\xi}^{\prime}$, respectively [31, 52, 53]. For a given spatial pattern $p(\mathbf{x},t)$, the spatial moments are defined as
$Z_{1}(t)$ $=\frac{1}{A}\int_{\Omega}p(\mathbf{x},t)\mathrm{d}\mathbf{x}$ (2.4a)
$Z_{2}(\bm{\xi},t)$ $=\frac{1}{A}\int_{\Omega}p(\mathbf{x},t)\left(p(\mathbf{x}+\bm{\xi},t)-\delta(\bm{\xi})\right)\mathrm{d}\mathbf{x}$ (2.4b)
$Z_{3}(\bm{\xi},\bm{\xi}^{\prime},t)$ $=\frac{1}{A}\int_{\Omega}p(\mathbf{x},t)\left(p(\mathbf{x}+\bm{\xi},t)-\delta(\bm{\xi})\right)\times$ (2.4c)
$\hskip 42.67912pt\left(p(\mathbf{x}+\bm{\xi}^{\prime},t)-\delta(\bm{\xi}^{\prime})-\delta(\bm{\xi}^{\prime}-\bm{\xi})\right)\mathrm{d}\mathbf{x}\,,$
in which the delta funtions in Eq. (2.4b) and Eq. (2.4c) prevent counting pairs and triplets with repeated organisms. Notice that we can obtain the radial correlation function from the definition of the second spatial moment (2.4b) as
[…]
To derive the dynamical equations for the spatial moments, we first need to compute the continuum analogue of the discrete neighbor-dependent death rate in Eq. (2.2). According to the definition in Eq. (2.2), the contribution of a neighboring individual located at a displacement $\bm{\xi}$, to the death rate of a focal plant is given by the competition kernel, $\phi_{C}(\bm{\xi})$. Now, the probability of an individual having a neighbor at a displacement, $\bm{\xi}$, is given by $Z_{2}(\bm{\xi},t)/Z_{1}(t)$ (see [53, 54] for detailed derivations). Consequently, the expected per-capita death rate can be obtained by multiplying $Z_{2}(\bm{\xi},t)/Z_{1}(t)$ by $g\phi_{C}(\bm{\xi})$ and integrating over all possible displacements as,
[…]
Since birth events are assumed to be neighbor-independent in the IBM [Eq. (2.1)], the expected per-capita birth rate is simply $B_{1}(t)=b$.
The dynamics of the first spatial moment are entirely governed by the balance between birth and death events because these two are the only processes influencing population size in our model. Therefore, the time evolution of the first spatial moment is:
$\frac{d}{dt}Z_{1}(t)=\Big{(}b-D_{1}(t)\Big{)}\,Z_{1}(t).$ (2.7)
Note that the dynamics of the first moment, in turn, are governed by the second moment through the death rate term in Eq. (2.6).
Next, we derive the dynamics of the second spatial moment. To this end, we first compute the death rate, $D_{2}(\bm{\xi},t)$, for a plant that is paired with another at a displacement $\bm{\xi}$. Remember that the neighbor-dependent component of the death rate for a single plant, $D_{1}(t)$, was conditioned on the presence of another plant at a displacement $\bm{\xi}$. In a similar fashion, the neighbor-dependent component of $D_{2}(\bm{\xi},t)$, is conditioned on the presence of a third plant at a displacement $\bm{\xi}^{\prime}$. This conditional probability of having another plant located at a displacement $\bm{\xi}^{\prime}$ is $Z_{3}(\bm{\xi},\bm{\xi}^{\prime},t)/Z_{2}(\bm{\xi},t)$. Following the same procedure as in Eq. (2.6), the expected death rate of a plant in pair with another at a separation displacement $\bm{\xi}$ can be computed by multiplying $Z_{3}(\bm{\xi},\bm{\xi}^{\prime},t)/Z_{2}(\bm{\xi},t)$ by $g\phi_{C}(\bm{\xi})$ and integrating over all possible displacements as follows:
[…]
Note that the additional factor of $g\phi_{C}(\bm{\xi})$ in the third term of Eq. (2.8) accounts for the direct influence of the individual at displacement $\bm{\xi}$.
We compute the dynamics of the second spatial moment by examining how the density of pairs at displacement $\bm{\xi}$ changes, balancing pair loss through death and creation through births. An existing pair of individuals separated by a displacement $\bm{\xi}$ can be lost if either one of the individuals dies, which occurs at a rate $D_{2}(\bm{\xi},t)$. Conversely, a new pair separated by $\bm{\xi}$ can be generated in two ways. First, if a pair already exists at displacement $\bm{\xi}+\bm{\xi}^{\prime}$, one of these individuals can disperse a daughter organism at a displacement $\bm{\xi}^{\prime}$, forming a new pair with displacement $\bm{\xi}$. This occurs at a rate $b\phi_{D}(\bm{\xi}^{\prime})$. Second, a new pair at displacement $\bm{\xi}$ can be created when a single individual disperses an offspring at displacement $-\bm{\xi}$, occurring at a rate $b\phi_{D}(-\bm{\xi})$. Combining these possibilities, we write the time evolution of the second spatial moment as,
[…]
The dynamics of the second moment still depend on the third moment through the definition of $D_{2}$ in (2.8), requiring a moment closure approximation to completely solve Eq. (2.9). Here, we employ the power-2 asymmetric closure:
$Z_{3}(\bm{\xi},\bm{\xi}^{\prime},t)=\frac{\Big{(}4Z_{2}(\bm{\xi},t)Z_{2}(\bm{\xi}^{\prime},t)+Z_{2}(\bm{\xi},t)Z_{2}(\bm{\xi}^{\prime}-\bm{\xi},t)+Z_{2}(\bm{\xi}^{\prime},t)Z_{2}(\bm{\xi}^{\prime}-\bm{\xi},t)-Z_{1}^{4}(t)\Big{)}}{5Z_{1}(t)},$ (2.10)
which approximates the third moment in terms of the first two spatial moments [55, 52].
2.3 Field equation for the biomass density field
Additionally, we can develop a coarse-grained equation for the biomass density field. This approximation is particularly useful when the spatial logistic model exhibits periodic spatial patterns. To derive this equation, we first discretize the two-dimensional square domain of lateral length $L$ used in the IBM definition into $m$ square cells of lateral length $\mathrm{d}x=L/m$. Each cell can be labeled with a pair of indices, $(i,j)$ where $i,j=1,...,m$, represent the cell coordinates in the discrete grid.
In this discretized space, the system’s state is given by the number of individuals in each grid cell $\mathbf{\Omega}\equiv\{N_{11},N_{12},...,N_{mm}\}$. At a given time $t$, the population is found in one of these states with a probability $P(\mathbf{\Omega},t)$ that can be used to compute the expected number of individuals in a focal cell
$\langle N_{ij}\rangle=\sum_{\mathbf{\Omega}}N_{ij}P(\mathbf{\Omega},t),$ (2.11)
where the average is over an ensemble of independent realizations of the individual-based stochastic dynamics. We can formally define the dynamics of this expected occupation number as
$\frac{d\langle N_{ij}\rangle}{dt}=\sum_{\mathbf{\Omega}}N_{ij}\frac{\partial P(\mathbf{\Omega},t)}{\partial t},$ (2.12)
where the time derivative of $P(\mathbf{\Omega},t)$ will be given by a master equation of the form
$\frac{\partial P(\mathbf{\Omega},t)}{\partial t}=\sum_{\mathbf{\Omega}^{\prime}}{\cal{T}}(\mathbf{\Omega}^{\prime}\rightarrow\mathbf{\Omega})P(\mathbf{\Omega}^{\prime},t)-{\cal{T}}(\mathbf{\Omega}\rightarrow\mathbf{\Omega}^{\prime})P(\mathbf{\Omega},t)$ (2.13)
where the sum in $\mathbf{\Omega}^{\prime}$ runs over all possible states and ${\cal{T}}$ are the transition rates between states and are given by the rules of the individual-based stochastic model. Because the probability of simultaneous births and deaths events is negligible, the IBM stochastic dynamics is a one step process [56]. In terms of the Master equation (2.13), this means that the transition rates $\cal{T}$ between pairs of states that differ in more than one particle are all zero. With this condition and using step operators [57], we can write the Master equation as
$\frac{\partial P(\mathbf{\Omega},t)}{\partial t}=\sum_{kl}\left(E_{kl}^{-}-1\right){\cal{T}}(\mathbf{\Omega}\rightarrow\mathbf{\Omega}_{kl}^{+})P(\mathbf{\Omega},t)+\left(E_{kl}^{+}-1\right){\cal{T}}(\mathbf{\Omega}\rightarrow\mathbf{\Omega}_{kl}^{-})P(\mathbf{\Omega},t),$ (2.14)
where we have defined the up-down step operators such that
$E_{ij}^{\pm}[f(\mathbf{\Omega})]=f[\left\{N_{11},...,N_{ij}\pm 1,...,N_{mm}\right\}]$ (2.15)
and the states $\mathbf{\Omega}_{kl}^{\pm}$ are identical to $\mathbf{\Omega}$ but have ±$\pm 1$ individual in the $kl$-th element. To fully describe the stochastic dynamics, we need to define the transition rates,
${\cal{T}}(\mathbf{\Omega}\rightarrow\mathbf{\Omega}_{kl}^{+})$ $=$ $\sum_{ij}b\,N_{ij}\phi_{\mathrm{D}}(\rvert\mathbf{x}_{ij}-\mathbf{x}_{kl}\rvert)\mathrm{d}x^{2},$ (2.16)
${\cal{T}}(\mathbf{\Omega}\rightarrow\mathbf{\Omega}_{kl}^{-})$ $=$ $N_{kl}\left(d+g\sum_{{ij}}N_{ij}\phi_{\mathrm{C}}(\rvert\mathbf{x}_{ij}-\mathbf{x}_{kl}\rvert)\right).$ (2.17)
Inserting the Master equation (2.14) with the transition rates in Eqs. (2.16)-(2.17) into the definition for the time derivative of the mean number of particles in each spatial cell, Eq. (2.12), we obtain
$\frac{d\langle N_{ij}\rangle}{dt}=b\sum_{kl}\phi_{D}(\rvert\mathbf{x}-\mathbf{x}^{\prime}\rvert)\mathrm{d}x^{2}\langle N_{kl}\rangle-d\langle N_{ij}\rangle-g\sum_{kl}\phi_{C}(\rvert\mathbf{x}-\mathbf{x}^{\prime}\rvert)\langle N_{ij}N_{kl}\rangle,$ (2.18)
where we have used the constraint that $\langle N_{ij}\rangle\geq 0$ $\forall ij$ and thus the probability of states with a negative number of particles, and the transition rates into them, must be zero. Finally, defining a scaled density of individuals, $n(\mathbf{x}_{ij})=N_{ij}g\,/\,b\mathrm{d}x^{2}$, and time $\tau=bt$ we obtain a partial differential equation for a population density field
$\frac{\partial\langle n\rangle}{\partial\tau}(\mathbf{x},\tau)=\int\phi_{D}(\rvert\mathbf{x}-\mathbf{x}^{\prime}\rvert)\langle n(\mathbf{x^{\prime}})\rangle\mathrm{d}\mathbf{x^{\prime}}-\frac{d}{b}\langle n(\mathbf{x})\rangle-\int\phi_{C}(\rvert\mathbf{x}-\mathbf{x}^{\prime}\rvert)\langle n(\mathbf{x})n(\mathbf{x^{\prime}})\rangle\mathrm{d}\mathbf{x^{\prime}},$ (2.19)
where we have omitted the time dependence on the right side. Notice that, similarly to what we obtained in the spatial moment dynamics equations in Section 2.2, Eq. (2.19) does not provide a closed description for the dynamics of the population density because it also includes its spatial covariance. We could write an equation for the dynamics of this spatial covariance and perform a moment closure approximation, as we did with the spatial moment dynamics equations. This approach is now much more complicated because the plant density is a field rather than a scalar quantity obtained from a spatial average. Therefore, we explore two alternative ways of treating these spatial correlations.
2.3.1 Mean-field approximation and the nonlocal Fisher KPP equation
The simplest and most common approximation to treat spatial correlations in Eq. (2.19) is to neglect the spatial covariance in the density field,
$\text{Cov}(\mathbf{x},\mathbf{x}^{\prime})=\langle n(\mathbf{x})n(\mathbf{x^{\prime}})\rangle-\langle n(\mathbf{x})\rangle\langle n(\mathbf{x^{\prime}})\rangle=0$ (2.20)
which implies that
$\langle n(\mathbf{x})n(\mathbf{x^{\prime}})\rangle=\langle n(\mathbf{x})\rangle\langle n(\mathbf{x^{\prime}})\rangle.$ (2.21)
Replacing Eq. (2.21) into (2.19), we obtain
$\frac{\partial\langle n\rangle}{\partial\tau}(\mathbf{x},\tau)=\int\phi_{D}(\rvert\mathbf{x}-\mathbf{x}^{\prime}\rvert)\langle n(\mathbf{x^{\prime}})\rangle\mathrm{d}\mathbf{x^{\prime}}-\frac{d}{b}\langle n(\mathbf{x})\rangle-\langle n(\mathbf{x})\rangle\int\phi_{C}(\rvert\mathbf{x}-\mathbf{x}^{\prime}\rvert)\langle n(\mathbf{x^{\prime}})\rangle\mathrm{d}\mathbf{x^{\prime}},$ (2.22)
that is an extension of the Fisher-Kolmogorov equation with nonlocalities in the competition and the growth terms. Defining a change of variables $\mathbf{y}\equiv\mathbf{x}-\mathbf{x}^{\prime}$ in the first integral of Eq. (2.22) and expanding the density $\langle n(\mathbf{x}-\mathbf{y})\rangle$ about $\mathbf{x}$ as a Taylor expansion up to second order, Eq. (2.22) reduces to
$\frac{\partial\langle n\rangle}{\partial\tau}(\mathbf{x},\tau)=\langle n(\mathbf{x})\rangle-\frac{d}{b}\langle n(\mathbf{x})\rangle-\langle n(\mathbf{x})\rangle\int\phi_{C}(\rvert\mathbf{x}-\mathbf{x}^{\prime}\rvert)\langle n(\mathbf{x^{\prime}})\rangle\mathrm{d}\mathbf{x^{\prime}}+\frac{\omega_{2}}{2}\frac{\partial^{2}}{\partial x^{2}}\langle n(\mathbf{x})\rangle,$ (2.23)
which is the nonlocal Fisher KPP equation with a diffusion coefficient defined by the reproduction rate (here absorbed in the scaled time) and the second moment of the dispersal kernel [58]. However, because we are interested in how long-range dispersal influences spatial patterns, we will retain the nonlocal dispersal term.
2.3.2 Beyond mean-field: introducing the effect of spatial covariances
Because Eq. (2.19) is isotropic and homogeneous, we propose an ansatz for the covariance of the form
$\text{Cov}(\mathbf{x},\mathbf{x}^{\prime})=\langle n(\mathbf{x})\rangle\langle n(\mathbf{x^{\prime}})\rangle f(|\mathbf{x}-\mathbf{x^{\prime}}|),$ (2.24)
where $f(|\mathbf{x}-\mathbf{x^{\prime}}|)$ is such that its modulus decays with its argument. In particular, to account for the fact that the spatial correlations of the density field go to zero when two points are very far from each other, $f(r)\rightarrow 0$ as $r\rightarrow\infty$ (where $r\equiv|\mathbf{x}-\mathbf{x^{\prime}}|$). Inserting this ansatz into the definition of the covariance in Eq. (2.20), we can write the second moment of the density field as,
$\langle n(\mathbf{x})n(\mathbf{x^{\prime}})\rangle=\langle n(\mathbf{x})\rangle\langle n(\mathbf{x^{\prime}})\rangle\left[1+f(|\mathbf{x}-\mathbf{x^{\prime}}|)\right],$ (2.25)
that we can insert into Eq. (2.19) to obtain
$\frac{\partial\langle n\rangle}{\partial\tau}(\mathbf{x},\tau)=\int\phi_{D}(|\mathbf{x}-\mathbf{x^{\prime}}|)\langle n(\mathbf{x^{\prime}})\rangle\mathrm{d}\mathbf{x^{\prime}}-\frac{d}{b}\langle n(\mathbf{x})\rangle-\langle n(\mathbf{x})\rangle\int\phi_{C}(|\mathbf{x}-\mathbf{x^{\prime}}|)\left(1+f(|\mathbf{x}-\mathbf{x^{\prime}}|)\right)\langle n(\mathbf{x^{\prime}})\rangle\mathrm{d}\mathbf{x^{\prime}}.$ (2.26)
Provided that we know the form of the function $f$, this equation fully describes the dynamics of the biomass density field, replacing the spatial correlations $\langle n(\mathbf{x})n(\mathbf{x^{\prime}})\rangle$ by an effective competition kernel that includes the ansatz function $f$.
To obtain an expression for this function $f$, we average Eq. (2.24) over space
$\int\text{Cov}(\mathbf{x},\mathbf{x+x^{\prime}})\mathrm{d}\mathbf{x}=f(|\mathbf{x^{\prime}}|)\int\langle n(\mathbf{x})\rangle\langle n(\mathbf{x}+\mathbf{x^{\prime}})\rangle\mathrm{d}\mathbf{x},$ (2.27)
from where we can write $f$ as a function of different features of the spatial pattern of population density,
$f(|\mathbf{x^{\prime}}|)=\frac{\int\text{Cov}(\mathbf{x},\mathbf{x+x^{\prime}})\mathrm{d}\mathbf{x}}{\int\langle n(\mathbf{x})\rangle\langle n(\mathbf{x}+\mathbf{x^{\prime}})\rangle\mathrm{d}\mathbf{x}}.$ (2.28)
All the quantities on the right side of Eq. (2.28) can be computed from IBM simulations. Alternatively, these quantities can also be obtained using SMD equations, which is computationally more efficient and makes it possible to have a closed dynamical equation for the vegetation density field without performing IBM simulations. Using the definitions of the spatial moments, we can rewrite (2.28) as,
$f(|\mathbf{x}-\mathbf{x^{\prime}}|)=\frac{Z_{2}(|\mathbf{x}-\mathbf{x^{\prime}}|)}{Z_{1}^{2}}-1,$ (2.29)
where we have used the pattern isotropy to write the dependence of the second spatial moment on the distance between two points explicitly. Using this relationship, we can describe the dynamics of the biomass density field using a system of two coupled equations: we use the equation for the second spatial moment to determine the density correlations and the function $f$, which we then insert in Eq. (2.26), to obtain the density field. In the next sections, we compare the accuracy of all these approximations to reproduce the total biomass density and pattern structure obtained in simulations of the stochastic IBM.
3 Results and Discussion
3.1 Spatial patterns in the discrete model
We first review the behavior of the spatially explicit spatial logistic model presented in Section 2.1 using a Gaussian competition kernel with standard deviation $r_{c}$ [36, 30] and a Gaussian dispersal kernel with standard deviation $s$. Holding the demographic parameters, $b$, $d$, and $g$ constant, the spatial correlations in the distribution of organisms are controlled by the dispersal and competition range.
Figure 1: Spatial moment dynamics results using a Gaussian competition kernel. A) Radial correlation function at displacement zero, $C(0)$ [36]. B) Total biomass density normalized by the mean-field expectation, $n_{\mathrm{MF}}=(b-d)/g$, as a function of the dispersal and competition ranges, $s$ and $r_{c}$ respectively. C) Radial correlation function for three combinations of $s$ and $r_{c}$ representative of each type of spatial pattern. Each curve is labeled with a symbol that indicates the values of $s$ and $r_{c}$ in (B). $b$, $d$, and $g$ with $d/b=0.1$ and $g=1$
Previous work has identified four possible types of spatial patterns depending on how these two scales compare to each other [36]. When both ranges are large, the dynamics are effectively non-spatial, and the IBM converges to the classical logistic model. However, when the dispersal scale decreases, while keeping competition long-ranged, individuals tend to clump, which leads to smaller population sizes than in the non-spatial case. Likewise, when the competition range decreases while keeping dispersal long-range, individuals tend to overdisperse, which increases the total population size. Finally, when both competition and dispersal are short-range, newborns cannot escape from competition with their ancestors and the population goes extinct. We recover these results integrating the SMD approximation in Eqs. (2.7)-(2.9) with the moment closure in Eq. (2.10) (Fig. 1a, b). In none of the scenarios discussed above, however, the spatial patterns are periodic (see radial correlation functions in Fig. 1c). Therefore, any of the continuous approximations we derived for the plant biomass density will return a homogeneous solution.
To investigate whether regular patterns exist in our model, we solved the SMD equations using a top-hat kernel as defined in Eq. (2.3) and compared these predictions with results from simulations of the discrete IBM. The IBM simulations show the emergence of regular patterns when the dispersal range is short and the competition range long (Fig. 2). This result is in good agreement with previous work using various nonlocal extensions of the Fisher KPP equation with top-hat competition kernels [46, 51]. We find that the SMD equations accurately predict the value of the radial correlation function at pairwise distance zero (top panels in Fig. 3b, c), which is a proxy for local crowding. The SMD equations, however, predict less accurately the population density. Specifically, when the competition range is large, and the dispersal range is short, the SMD equations underestimate the total population size resulting from the IBM simulations (central panels in Fig. 3a, b).
Figure 2: A) Long-time periodic pattern snapshot with dispersal range $s=0.15$ and competition range $r_{c}=2.2$ (inside the region of high aggregation in the $C(0)$ heatmaps in Fig. 3). B) Fourier transform of the pattern snapshot.
This discrepancy in the total population density comes from the inaccurate prediction of the full radial correlation function provided by the SMD equations. In particular, SMD does not fully capture the amplitude of the oscillations in the radial correlation functions measured from IBM simulations (rightmost panel in the bottom row of Fig. 3b, c). This amplitude, especially when the radial correlation function becomes negative, is critical to determine how larger the population density grows when regular patterns form. More negative values of $C(r)$ indicate that the regions between clusters are less populated, which favors the growth of more individuals within each aggregate [47].
Note that the moment closure method employed here to approximate the third moment in terms of the first two moments inherently relies on two key assumptions. First, the third moment must be positive, as it represents triplet density and is calculated from the product of three non-negative densities (Eq. (2.4c)). Therefore, the closure approximation must also remain non-negative. The second assumption is the dynamical invariance under relabeling, implying that relabeling identical individuals should not alter the system dynamics [55]. Different closure methods balance these assumptions, with the power-2 closure (used in our computations) offering an optimal trade-off between them. However, in certain parameter regimes, these assumptions may not hold, leading to notable discrepancies between IBM and SMD [59, 55] as we observed in the radial correlation function of the periodic patterns (diamond panels in Fig. 3). Additionally, truncating the hierarchy of spatial moments at the second order may limit model accuracy for spatial structures with significant amounts of information in spatial moments at higher orders.
Figure 3: Discrete model analysis using a top-hat competition kernel. A) Snapshots of spatial patterns. From left to right: segregated, uniform, and regular aggregated. Notice that to facilitate the visualization of the pattern spatial structure, the aggregated and uniform patterns are shown only in a small region of the spatial domain. B) From top to bottom, SMD predictions for the radial correlation function at zero displacement, $C(0)$, the total density of plants normalized by the homogenous expectation, and the radial correlation function for the three combinations of $s$ and $r_{c}$ used in (A) and indicated by the coordinates of the red symbols in the $C(0)$ panel. and carrying capacity derived from spatial moment dynamics. C) Same properties as B) but computed from direct simulations of the stochastic IBM.
3.2 Pattern formation in the field equation
To test the accuracy of the field equation in predicting the behavior of the individual-based simulations, we performed a linear stability analysis around the non-trivial homogenous solution to obtain the onset of pattern formation. Doing this analysis on the coupled system of equations for the second spatial moment $Z_{2}$ and the density field $\langle n\rangle$ is very challenging because the equation for $Z_{2}$ is inhomogeneous (it depends explicitly on space) and thus does not have any homogeneous equilibrium apart from the vanishing density field for which $Z_{2}=0$ too. Therefore, instead of working with the full system of coupled equations, we use the results from the SMD numerical simulation and assume that $Z_{2}=(1+f)Z_{1}^{2}$ is a fixed point of $\dot{Z_{2}}=0$. Inserting this solution into the equation for $\langle n\rangle$, we obtain the density equilibrium
$\langle n\rangle_{h}=\frac{1-\frac{d}{b}}{1+\int\phi_{C}(|\mathbf{x^{\prime}}|)f(|\mathbf{x^{\prime}}|)\mathrm{d}\mathbf{x^{\prime}}}.$ (3.1)
We can perform a linear stability analysis of this equilibrium for small spatiotemporal perturbations, considering that $Z_{2}$ (and hence $f$) is given. This calculation provides the linear growth rate of the density Fourier modes
$\lambda(\mathbf{k})=F[{\phi}_{D}](\mathbf{k})-\langle n\rangle_{h}\,F\left[\phi_{C}(1+f)\right](\mathbf{k})-\left(\frac{d}{b}+\langle n\rangle_{h}\int\phi_{C}(|\mathbf{x^{\prime}}|)f(|\mathbf{x^{\prime}}|)\mathrm{d}\mathbf{x^{\prime}}\right),$ (3.2)
where $F[]$ indicates a Fourier transform operation. The onset of pattern formation is defined by the parameter values such that $\lambda(\mathbf{k}_{c})=0$, where $\mathbf{k}_{c}$ is the non-zero Fourier mode at which the perturbation growth rate is maximum. Therefore, we can define the region where patterns might exist by scanning the $s,r_{c}$ parameter space and testing for the pattern onset condition.
To quantify how accurately this linear stability analysis predicts the region of the IBM in which regular patterns form, we need to measure regularity in the stochastic patterns obtained from the discrete simulations. We introduce two different Fourier-based metrics, a first one based on comparing the distribution of Fourier spectra generated by simulated and random patterns, and a second one based on how strongly the pattern wavenumber (maximum of the pattern power spectrum) contributes to the total power of the pattern spectrum (see Appendix A for details on how to compute these metrics). As expected, both metrics identify that regular patterns form when the dispersal range is short and the competition range is large (bottom right corner in Fig. 4). When comparing these results with the predictions of the field equation, we find that the mean-field approximation (gray line in Fig. 4) underestimates the region of the $s,r_{c}$ parameter space where periodic patterns exist. However, accounting for spatial correlations, either using the SMD equations or IBM simulations, provides a much better agreement with the results of the discrete model (cyan symbols and red dashed line in Fig. 4). Additionally, the agreement between the corrected PDE and the IBM simulations is much better when regularity in the discrete patterns is measured using the relative amplitude of the pattern wavenumber. This metric provides a much more robust measurement of regularity because it does not rely on any statistical modeling assumption. The first metric we introduced, however, quantifies regularity as a statistically significant difference between random patterns and those resulting from the IBM simulation, which makes the definition of regularity sensitive to both the number of pattern realizations we used to construct the statistical ensembles and the percentiles used to define the difference between power spectra distributions. As a result, the transition from regular to non-regular patterns, indicated by the white region in Fig. 4a, will vary depending on how we choose these quantities.
Figure 4: Validation of how the different PDE approximations predict regular patterns. A) Regularity is quantified by comparing IBM simulated and random patterns. B) Regularity is measured by the power of the pattern wavelength normalized by the total power (see App. A for details on each metric). Curves in the panels correspond to each approximation in the treatment of the spatial correlations (see legend in panel (B).
4 Conclusions
We investigated the spatial patterns that form in a spatially explicit IBM logistic model with nonlocal dispersal and intraspecific competition using both a Gaussian and top-hat competition kernel. Combining numerical simulations of the discrete model and a spatial moment dynamics approximation, we found that, for top-hat competition kernels, the population might exhibit periodic patterns with aggregates of individuals forming a hexagonal lattice [46]. This type of pattern is not present when long-range competition is modeled using a Gaussian kernel [36]. Interestingly, the emergence of these regular patterns also affects the system productivity, allowing for larger population sizes even in conditions where local individual crowding is high [51].
Next, we looked at how accurately a field equation for the biomass density, obtained by upscaling the individual-based dynamics, reproduced the existence of regular patterns. We found that the standard mean-field approximation, which neglects spatial covariates in the density field, does not fully capture the region of the parameter space with regular patterns. To improve this prediction, we developed a novel approximation that uses spatial-moment dynamics equations to introduce the effect of spatial covariates in the density field. This corrected density equation accurately reproduces the region with regular patterns, offering a new avenue to study pattern formation in sessile organisms using continuous, density-based models beyond mean-field approximations.
Acknowledgments
This work was partially funded by the Center of Advanced Systems Understanding (CASUS), which is financed by Germany’s Federal Ministry of Education and Research (BMBF) and by the Saxon Ministry for Science, Culture and Tourism (SMWK) with tax funds on the basis of the budget approved by the Saxon State Parliament. RMG was supported by FAPESP through grant ICTP-SAIFR 2021/14335-0. RM was supported by CNPq through grant 140096/2021-3 and CAPES - Finance Code 001.
Appendices
App. A Computation of the regularity metrics
We introduced two different Fourier-based metrics to quantify regularity in the IBM spatial patterns:
[…]
  • The second metric quantifies the periodicity in the IBM simulated patterns measuring the amplitude of the characteristic wavenumber (the wavenumber at which the radial power spectrum is maximum) to the total power, computed as the integral of the radial power spectrum across the entire range of wavenumber.
[…]